# Globalization of perturbative Chern-Simons theory on the moduli space of flat connections in the BV formalism

### v2

@@ Abstract.
K. W. was supported by the ERC-SyG project, Recursive and Exact New Quantum Theory (ReNewQuantum) with funding from the European Research Council under the European Union‘s Horizon 2020 research and innovation programme, grant agreement no. 810573, and by a Sapere Aude Grant from the Independent Research Fund Denmark, grant ID 10.46540/4251-00015B
@@ 1.1.2. Mathematical definition of the perturbative partition function
  • $\tau(A_{0})$ is the Ray-Singer torsion. For $A_{0}$ non-acyclic, rather than being a number, it is an element of the determinant line of the cohomology $H^{\bullet}_{A_{0}}$.
@@ 1.1.3. Aside: BV pushforward perspective
Also, the property (8) suggests that under a deformation of Riemannian metric on $M$, $Z$ should change by a $\Delta^{\prime}$-exact term.
@@ 1.1.5. The global partition function
    (15) $Z^{\mathrm{glob}}=\left(\sum_{k\geq 0}\frac{(i\hbar)^{k}}{k!}\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A_{0}]}\right\rangle^{k}\widetilde{Z}\right)\Bigg|_{\mathsf{a}^{1}=[\delta A_{0}]=0}.$
@@ 1.2.2. Main result 1
    (17) $\det(B^{\vee})\circ Z_{\phi(A_{0},\alpha)}(B(\mathsf{a}))=Z_{A_{0}}(\alpha+\mathsf{a})+i\hbar\Delta_{\mathsf{a}}R(A_{0},\alpha;\mathsf{a})$
@@
    • $\Delta_{\mathsf{a}}$ is the BV Laplacian on formal half-densities on the fiber of $\mathbb{T}\mathcal{M}^{\prime}$ over $A_{0}$;
@@
    (18) $\nabla^{G}Z=i\hbar\Delta_{\mathsf{a}}R_{1}$
@@
    (23) $Z_{A,A_{1}^{\prime}}(\mathsf{a})=Z_{A,A^{\prime}_{0}}(\mathsf{a})+i\hbar\Delta_{\mathsf{a}}R(A,A^{\prime}_{0},A_{1}^{\prime},\mathsf{a}).$
@@
    (24) $\delta_{A^{\prime}}Z_{A,A^{\prime}}(\mathsf{a})=i\hbar\Delta_{\mathsf{a}}R_{\delta A^{\prime}}(A,A^{\prime},\mathsf{a}),$
@@ 1.4. Metric (in)dependence of the global partition function – main result 3
    (26) $\delta_{g}Z^{\mathrm{ren}}_{A_{0}}(\mathsf{a})=i\hbar\Delta_{\mathsf{a}}R_{\delta g}.$
@@
(28) $\frac{d}{dt}\bigg|_{t=0}Z_{t}^{\mathrm{glob,ren}}=i\hbar\Delta_{\mathcal{M}^{\prime}}(R^{\mathrm{glob}}),$
@@
(29) $\frac{d}{dt}\bigg|_{t=0}Z_{t}^{\mathrm{glob,ren}}=i\hbar\,d_{\mathcal{M}^{\prime}}R^{\mathrm{glob}}.$
@@ 1.7. Comparison to literature and historical remarks
The problem of studying the perturbative (or semiclassical) behavior of the Chern-Simons partition function around non-acyclic flat connections, where the path integral has zero modes, was already observed in Witten’s seminal paper on the subject [Wit89, p.361]. Axelrod and Singer studied the perturbative theory around acylic flat connections in detail [AS91], [AS94] but already comment that the assumption on acyclicity should be removed, and state (without proof) that the partition function changes by a total divergence when changing the Riemannian metric. They also identify the problem of defining the integral over the moduli space and proving that it is finite, as well as potential anomalies. Axelrod has a later preprint on the subject [Axe95], where he develops the theory of oscillatory integrals of Morse-Bott functions and announces some theorems on their application to Chern-Simons theory, but without proof. Our work in this paper is independent from this preprint [^25] In fact, we only learned about its existence shortly before completion of this paper. and draws on a different background - BV pushforwards. The main body of the literature on perturbative Chern-Simons theory turned to the study of (rational) homology 3-spheres, where one can treat the problem of zero modes either by puncturing [Kon93a], [KT99], [Les02] (resulting in the Kontsevich-Kuperberg-Thurston-Lescop or KKTL invariant) or by introducing extra vertices as in the works of Bott and Cattaneo [BC98], [BC99] to cancel the effect of zero modes. [^26] Here one should mention the recent paper [CS21] filling a gap in the construction of Bott and Cattaneo. Cattaneo later showed those constructions agree [Cat99]. Another line of research focused on extracting perturbative invariants of 3-manifolds from the Kontsevich integral [Kon93], such as the Aarhus integral [BGRT02], [BGRT02a] and the LMO invariant [LMO98]. [^27] It is known that the Aarhus integral and the LMO invariant are equivalent [BGRT04]. It is conjectured that the KKTL invariant and the LMO are equal, but this is known only up to degree 2 for integral homology spheres (accredited to C. Lescop in private communication of K.W. with G. Massuyeau). A full definition of the perturbative Chern-Simons partition function at non-acyclic flat connections only appeared with the introduction of the BV formalism to the problem and the works of Cattaneo and the first author [CM08] (see also [Mne19], [Wer22]) and simultaneously Iacovino [Iac08].
@@ 2.1. Smooth points
the cohomology of $d_{A_{0}}$. By homotopy transfer of $L_{\infty}$-algebras, $H^{\bullet}_{A_{0}}$ is turned into a minimal $L_{\infty}$-algebra endowed with induced operations $\{l^{\prime}_{n,A_{0}}\}_{n\geq 2}$. A choice of SDR data282828“Strong Deformation Retraction data” [GL89], also known in the literature under the names “contraction” [EL53], “homotopy equivalence data” [Cra04], “induction data,” “$(i,p,K)$ triple.” $r_{A_{0}}=(i_{A_{0}},p_{A_{0}},K_{A_{0}})$ of $\Omega^{\bullet}(M,\mathfrak{g})$ onto $H^{\bullet}_{A_{0}}$ provides us with explicit representatives of these operations (see Appendix A for our conventions on SDR data). Denote $T_{n}$ the set of isomorphism classes of binary rooted trees with $n$ leaves — here we think of leaves and the root as half-edges emanating from internal vertices. To $T\in T_{n}$ we can associate an $n$-ary operation $\lambda_{T}\colon\wedge^{n}H^{\bullet}_{A_{0}}\to H^{\bullet}_{A_{0}}$ of degree 2n2-n as follows: To the $n$ leaves we assign the map $i_{A_{0}}$, to internal vertices we assign the map $l_{2}$, to internal edges we assign the map $K_{A_{0}}$, to the root we assign the map $p_{A_{0}}$ (see Figure 2); then we skew-symmetrize over the permutations of $n$ inputs on the $n$ leaves.
@@
The purpose of this subsection is to prove that both the space of all flat connections and the moduli space of flat connections are smooth manifolds close to a smooth point $A_{0}$, resp. its class in the moduli space $A_{0}]$. It follows that $\mathrm{FC}^{\mathrm{sm}}$ and $\mathcal{M}^{\mathrm{sm}}=\pi(\mathrm{FC}^{\mathrm{sm}})$ are smooth manifolds. [^29] When going to the moduli space, we will only prove it for the subset of smooth irreducible flat connections $\mathcal{M}^{\prime}=\pi(\mathrm{FC}^{\prime})$, i.e. flat connections $A_{0}$ for which $H^{0}_{A_{0}}=0.$ In the infinite-dimensional case, we work in the Banach setting and assume the following:
Assumption 2.2 (Banach boundedness).
There is a Banach norm $||\cdot||_{B}$ on $\Omega^{\bullet}$ such that $r_{A_{0}}=(i_{A_{0}},p_{A_{0}},K_{A_{0}})$ are bounded linear maps with respect to $||\cdot||_{B}$ and some norm on $H^{\bullet}_{{A_{0}}}$.
In our examples, this Banach norm will be a Sobolev norm. We denote by $\Omega^{\bullet}_{B},\Omega^{k}_{B},\mathrm{FC}_{B},\ldots$ the completion of those space with respect to $||\cdot||_{B}$. We shall prove the following theorem:
@@
The set of smooth points $\mathrm{FC}^{\mathrm{sm}}\subset\mathrm{FC}$ has the structure of a Banach manifold. For every smooth point $A_{0}$, there is a neighborhood $V_{A_{0}}$ of $A_{0}\in\mathrm{FC}_{B}$ modeled on a neighborhood $U_{A_{0}}$ of $0\in(\Omega^{1}_{A_{0}-\mathrm{cl},B},||\cdot||_{B})$. Given SDR data $r_{A_{0}}$ at $A_{0}$ satisfying Assumption 2.2, we have a chart
@@ 2.1.1. Formal deformations of flat connections
We recall the following elementary facts about flat connections. If $A_{t}\colon(-\epsilon,\epsilon)\to\Omega^{1}(M,\mathfrak{g})$ is a smooth curve of flat connections, then from differentiating $F_{A_{t}}=0$ we obtain $d_{A_{0}}\dot{A}_{0}=0$, i.e., tangent vectors at $A_{0}$ to curves of flat connections are $d_{A_{0}}$-closed 1-forms. If $\mathrm{g}_{t}\colon(-\epsilon,\epsilon)\to C^{\infty}(M,G)$ is a curve with $\mathrm{g}_{0}\equiv 1$ and $\dot{\mathrm{g}}_{0}=\gamma\in\Omega^{0}(M,\mathfrak{g})$, then the tangent vector at 0 to the curve of flat connections $A_{t}={}^{\mathrm{g}_{t}}A_{0}$ is $\dot{A}_{0}=-d_{A_{0}}\gamma$, i.e. the tangent vector to a curve along the gauge orbit of a flat connection is a $d_{A_{0}}$-exact 1-form. This is equivalent to saying that infinitesimal [^30] By “infinitesimal” we everywhere mean “first-order,” as opposed to formal deformations (of infinite order) discussed below. deformations of flat connections are $d_{A_{0}}$-closed 1-forms, and two such deformations are equivalent whenever they differ by an exact 1-form, i.e. equivalence classes of infinitesimal deformations are in 1-to-1 correspondence with the first twisted cohomology group $H^{1}_{{A_{0}}}$ (sometimes called the Zariski tangent space to the moduli space of flat connections at $A_{0}]$).
@@
The first two equations are satisfied by our assumptions. It is instructive to look at the third equation in detail. We see that we can solve it for $\alpha^{(2)}$ if and only $\frac{1}{2}[\alpha^{(1)},\alpha^{(1)}]$ is $d_{A_{0}}$-exact. Because $\alpha^{(1)}=\alpha$ is closed by assumption, and the bracket is compatible with the differential, $\frac{1}{2}[\alpha,\alpha]$ is always $d_{A_{0}}$ closed. It is exact if and only if $l_{2}^{\prime}([\alpha,\alpha])=0$. In this case, we can write down an explicit solution:

### v3

@@ Abstract.
P.M. acknowledges partial support of the Simons Collaboration on Global Categorical Symmetries, of the Simons Foundation Travel Support Grant, and of the FIM at ETH Zurich. P.M.’s research was (partly) supported by the NCCR SwissMAP, funded by the Swiss National Science Foundation. K. W. was supported by the ERC-SyG project, Recursive and Exact New Quantum Theory (ReNewQuantum) with funding from the European Research Council under the European Union’s Horizon 2020 research and innovation programme, grant agreement no. 810573, and by a Sapere Aude Grant from the Independent Research Fund Denmark, grant ID 10.46540/4251-00015B.
@@ 1.1.2. Mathematical definition of the perturbative partition function
  • $\tau(A_{0})$ is the Ray-Singer torsion. For $A_{0}$ non-acyclic, rather than being a number, it is an element of the determinant line of the cohomology $H^{\bullet}_{A_{0}}$, see Remark 1.1 below.
@@ 1.1.3. Aside: BV pushforward perspective
Remark 1.1.
The free Gaussian part of the BV pushforward (11) is the formal Gaussian integral of $e^{\frac{i}{\hbar}\int_{M}\mathrm{tr}\frac{1}{2}B\wedge d_{A_{0}}B}$ which is interpreted (as guided by finite-dimensional oscillatory Gaussian integrals) as
(12) $({\det}^{\mathrm{reg}}d_{A_{0}})^{\frac{1}{2}}e^{\frac{\pi i}{4}\mathrm{sign}^{\mathrm{reg}}(d_{A_{0}})},$}"
with ${\det}^{\mathrm{reg}}d_{A_{0}}$” the gauge-fixed (gauge-fixing can be thought of replacing $d_{A_{0}}$ with $(d^{*}_{A_{0}}d_{A_{0}})^{\frac{1}{2}}$ and then restricting it to coexact forms) and zeta-regularized determinant of $d_{A_{0}}$, a.k.a. the Ray-Singer torsion. We refer to [Sch78], [Wit89], [Mne14, Section 1.7], [CMR20] for details on Ray-Singer torsion $\tau(A_{0})$ (or Reidemeister torsion in the combinatorial setting of [CMR20]) and how it appears as a Gaussian integral with kinetic operator given by de Rham operator twisted by a flat connection. The gauge-fixed and zeta-regularized signature of the kinetic operator in (12) is understood as the Atiyah-Patodi-Singer eta-invariant $\psi(A_{0})$, see [Wit89].
Ray-Singer torsion is the regularized version of the general algebraic notion of torsion of a cochain complex $C^{\bullet}$ with inner product. This torsion $\tau(C^{\bullet})$ is an element of the determinant line $\mathrm{Det}\,C^{\bullet}\cong\mathrm{Det}\,H^{\bullet}$, with $H^{\bullet}$ the cohomology of $C^{\bullet}$, see [Mil66], [Mne14]. (In particular, if $C^{\bullet}$ is acyclic, the torsion is a number.) In the case of $C^{\bullet}$ being the de Rham complex twisted by a flat connection $A_{0}$, the torsion becomes the Ray-Singer torsion – an element of the determinant line of twisted de Rham cohomology $H^{\bullet}_{A_{0}}$, or, equivalently, a constant density on the graded space $H^{\bullet}_{A_{0}}[1]$. Thus, the square root of torsion participating in the perturbative Chern-Simons path integral (4) is a (constant) half-density on the space of zero-modes $H^{\bullet}_{A_{0}}[1]$. We refer also to [BZ92], [Bra03] for a discussion of Ray-Singer torsion for non-acyclic flat connections.
Note that the property (8) suggests that under a deformation of Riemannian metric on $M$, $Z$ should change by a $\Delta^{\prime}$-exact term.
@@ 1.1.5. The global partition function
    (16) $Z^{\mathrm{glob}}=\left(\sum_{k\geq 0}\frac{(-i\hbar)^{k}}{k!}\left\langle\frac{\partial}{\partial\mathsf{a}^{2}},\frac{\partial}{\partial[\delta A_{0}]}\right\rangle^{k}\widetilde{Z}\right)\Bigg|_{\mathsf{a}^{1}=[\delta A_{0}]=0}.$
@@ 1.2.2. Main result 1
    (18) $\det(B^{\vee})\circ Z_{\phi(A_{0},\alpha)}(B(\mathsf{a}))=Z_{A_{0}}(\alpha+\mathsf{a})-i\hbar\Delta_{\mathsf{a}}R(A_{0},\alpha;\mathsf{a})$
@@
    • $\Delta_{\mathsf{a}}=\left\langle\frac{\partial}{\partial\mathsf{a}^{1}},\frac{\partial}{\partial\mathsf{a}^{2}}\right\rangle$, with brackets meaning the Poincaré duality pairing, is the BV Laplacian on formal half-densities on the fiber of $\mathbb{T}\mathcal{M}^{\prime}$ over $A_{0}$;
@@
    (19) $\nabla^{G}Z=-i\hbar\Delta_{\mathsf{a}}R_{1}$
@@
    (24) $Z_{A,A_{1}^{\prime}}(\mathsf{a})=Z_{A,A^{\prime}_{0}}(\mathsf{a})-i\hbar\Delta_{\mathsf{a}}R(A,A^{\prime}_{0},A_{1}^{\prime},\mathsf{a}).$
@@
    (25) $\delta_{A^{\prime}}Z_{A,A^{\prime}}(\mathsf{a})=-i\hbar\Delta_{\mathsf{a}}R_{\delta A^{\prime}}(A,A^{\prime},\mathsf{a}),$
@@ 1.4. Metric (in)dependence of the global partition function – main result 3
    (27) $\delta_{g}Z^{\mathrm{ren}}_{A_{0}}(\mathsf{a})=-i\hbar\Delta_{\mathsf{a}}R_{\delta g}.$
@@
(29) $\frac{d}{dt}\bigg|_{t=0}Z_{t}^{\mathrm{glob,ren}}=-i\hbar\Delta_{\mathcal{M}^{\prime}}(R^{\mathrm{glob}}),$
@@
(30) $\frac{d}{dt}\bigg|_{t=0}Z_{t}^{\mathrm{glob,ren}}=-i\hbar\,d_{\mathcal{M}^{\prime}}R^{\mathrm{glob}}.$
@@ 1.7. Comparison to literature and historical remarks
The problem of studying the perturbative (or semiclassical) behavior of the Chern-Simons partition function around non-acyclic flat connections, where the path integral has zero modes, was already observed in Witten’s seminal paper on the subject [Wit89, p.361]. Axelrod and Singer studied the perturbative theory around acylic flat connections in detail [AS91], [AS94] but already comment that the assumption on acyclicity should be removed, and state (without proof) that the partition function changes by a total divergence when changing the Riemannian metric. They also identify the problem of defining the integral over the moduli space and proving that it is finite, as well as potential anomalies. Axelrod has a later text in conference proceedings on the subject [Axe95], where he develops the theory of oscillatory integrals of Morse-Bott functions and announces some theorems on their application to Chern-Simons theory, but without proof. Our work is independent from Axelrod’s paper [^25] In fact, we only learned about its existence shortly before completion of this paper. and draws on a different background - BV pushforwards. The main body of the literature on perturbative Chern-Simons theory turned to the study of (rational) homology 3-spheres, where one can treat the problem of zero modes either by puncturing [Kon93a], [KT99], [Les02] (resulting in the Kontsevich-Kuperberg-Thurston-Lescop or KKTL invariant) or by introducing extra vertices as in the works of Bott and Cattaneo [BC98], [BC99] to cancel the effect of zero modes. [^26] Here one should mention the recent paper [CS21] filling a gap in the construction of Bott and Cattaneo. Cattaneo later showed those constructions agree [Cat99]. Another line of research focused on extracting perturbative invariants of 3-manifolds from the Kontsevich integral [Kon93], such as the Aarhus integral [BGRT02], [BGRT02a] and the LMO invariant [LMO98]. [^27] It is known that the Aarhus integral and the LMO invariant are equivalent [BGRT04]. It is conjectured that the KKTL invariant and the LMO invariant are equal, but this is known only up to degree 2 for integral homology spheres (accredited to C. Lescop in private communication of K.W. with G. Massuyeau). A full definition of the perturbative Chern-Simons partition function at non-acyclic flat connections only appeared with the introduction of the BV formalism to the problem and the works of Cattaneo and the first author [CM08] (see also [Mne19], [Wer22]) and simultaneously Iacovino [Iac08].
@@
1.8. Acknowledgements
We thank Alberto S. Cattaneo for inspiring discussions. We also thank the referees for their helpful comments.
@@ 2.1. Smooth points
the cohomology of $d_{A_{0}}$. By homotopy transfer of $L_{\infty}$-algebras, $H^{\bullet}_{A_{0}}$ is turned into a minimal $L_{\infty}$-algebra endowed with induced operations $\{l^{\prime}_{n,A_{0}}\}_{n\geq 2}$. A choice of SDR data282828“Strong Deformation Retraction data” [GL89], also known in the literature under the names “contraction” [EL53], “homotopy equivalence data” [Cra04], “induction data,” “$(i,p,K)$ triple.” $r_{A_{0}}=(i_{A_{0}},p_{A_{0}},K_{A_{0}})$ of $\Omega^{\bullet}_{A_{0}}$ onto $H^{\bullet}_{A_{0}}$ provides us with explicit representatives of these operations (see Appendix A for our conventions on SDR data). Denote $T_{n}$ the set of isomorphism classes of binary rooted trees with $n$ leaves — here we think of leaves and the root as half-edges emanating from internal vertices. To $T\in T_{n}$ we can associate an $n$-ary operation $\lambda_{T}\colon\wedge^{n}H^{\bullet}_{A_{0}}\to H^{\bullet}_{A_{0}}$ of degree 2n2-n as follows: To the $n$ leaves we assign the map $i_{A_{0}}$, to internal vertices we assign the map $l_{2}$, to internal edges we assign the map $K_{A_{0}}$, to the root we assign the map $p_{A_{0}}$ (see Figure 2); then we skew-symmetrize over the permutations of $n$ inputs on the $n$ leaves.
@@
The purpose of this subsection is to prove that both the space of all flat connections and the moduli space of flat connections are smooth manifolds close to a smooth point $A_{0}$, resp. its class in the moduli space $A_{0}]$. It follows that $\mathrm{FC}^{\mathrm{sm}}$ and $\mathcal{M}^{\mathrm{sm}}=\pi(\mathrm{FC}^{\mathrm{sm}})$ are smooth manifolds. [^29] When going to the moduli space, we will only prove it for the subset of smooth irreducible flat connections $\mathcal{M}^{\prime}=\pi(\mathrm{FC}^{\prime})$, i.e. flat connections $A_{0}$ for which $H^{0}_{A_{0}}=0.$
Remark 2.2 (Topologies on $\Omega^{\bullet}$ and $\mathrm{FC}$).
The vector space $\Omega^{\bullet}$ carries the natural Fréchet topology of uniform convergence, with respect to which it is complete. Moreover, it also carries the coarser Banach topologies induced by the Sobolev norms [^30] This norm is defined by the $L^{p}$-norms of $\omega$ and covariant derivatives up to order $l$, with respect to a given Riemannian metric $g$. $||\cdot||_{l,p}$. The spaces $\mathrm{FC}^{\prime}\subset\mathrm{FC}^{\mathrm{sm}}\subset\mathrm{FC}\subset\Omega^{1}$ carry the corresponding subset topologies. With respect to the $||\cdot||_{l,p}$ norm none of those spaces is complete and we denote the completion by $\Omega^{\bullet}_{l,p},\mathrm{FC}_{l,p},\ldots$. We usually work with $p=2$ and denote $\Omega^{\bullet}_{l,2}=\Omega_{l}$. [^31] The completions in the Sobolev topologies can contain connections with singularities. However, when we say a connection is smooth, we mean that it is a smooth point in $\mathrm{FC}$.
Given the different topologies, one can make different assumptions on $K_{A_{0}}$.
Assumption 2.3 ($l$-Sobolev boundedness).
The chain homotopy $K_{A_{0}}\colon\Omega^{\bullet}\to\Omega^{\bullet-1}$ is bounded with respect to the Sobolev norm $||\cdot||_{l,2}$.
Actually, the main example considered in this text, the SDR coming from Hodge decomposition (Example 2.30) satisfies the following stronger assumption:
Assumption 2.4 (Fréchet continuity).
The chain homotopy $K_{A_{0}}\colon\Omega^{\bullet}\to\Omega^{\bullet-1}$ is continuous in the Fréchet topology.
The goal of this subsection is to prove the following theorems:
@@
For every $l\geq 2$, the completion of the set of smooth points $\mathrm{FC}^{\mathrm{sm}}_{l}\subset\mathrm{FC}_{l}$ has the structure of a Banach manifold. For every smooth point $A_{0}$, there is a neighborhood $V_{A_{0}}$ of $A_{0}\in\mathrm{FC}_{l}$ modeled on a neighborhood $U_{A_{0}}$ of $0\in(\Omega^{1}_{A_{0}-\mathrm{cl},l},||\cdot||_{l})$. Given SDR data $r_{A_{0}}$ at $A_{0}$ satisfying Assumption 2.3, we have a homeomorphism in the Banach topologies
@@
The chart $\widetilde{\varphi}$ is constructed using the chain homotopy $K_{A_{0}}$. If $K_{A_{0}}$ is Fréchet continuous, then $\widetilde{\varphi}_{A_{0}}$ is actually a chart for the Fréchet manifold $\mathrm{FC}^{\mathrm{sm}}$.
Theorem 2.6.
The set $\mathrm{FC}^{\mathrm{sm}}\subset\mathrm{FC}$ has the structure of a Fréchet manifold. If the chain homotopy $K_{A_{0}}$ satisfies Assumption 2.4, then the map (40), restricted to $U_{A_{0}}\cap\Omega^{1}_{d_{A_{0}}-\mathrm{cl}}$, is a Fréchet homeomorphism.
In the later sections of the text, we consider $\mathrm{FC}^{\prime}\subset\mathrm{FC}^{\mathrm{sm}}$, with this Fréchet manifold structure.
2.1.1. Outline of the proof
We briefly outline the strategy of the proof. The Banach case was inspired by [HWW].
  1. (1)
    We define the formal power series $\delta_{A_{0}}=\sum_{k\geq 0}\alpha^{(k)}$ as a sum over trees, and show that $A_{0}+\delta_{A_{0}}$ is flat for $\alpha^{(1)}$ harmonic – Proposition 2.7.
  2. (2)
    We show that $\widetilde{\delta}_{A_{0}}$ (defined using the sum over trees for all closed forms) converges in the Banach topology under Assumption 2.3 (Proposition 2.9) or in the Fréchet topology under Assumption 2.4 (Remark 2.10).
  3. (3)
    We define the Kuranishi map (60), show that it is invertible on $\Omega^{1}$ and a compositional inverse of $\delta_{A_{0}}$ on $\Omega^{1}$, and show that it maps Maurer-Cartan forms to closed forms (Proposition 2.12).
  4. (4)
    We establish the Banach smoothness of $\mathrm{FC}^{\mathrm{sm}}$ by means of the chart $\widetilde{\psi}$, see Proposition 2.13.
  5. (5)
    Now, one can use the inverse function theorem for $\widetilde{\kappa}$ on $\mathrm{FC}^{\mathrm{sm}}$ to conclude that $\widetilde{\kappa}$, restricted to the Maurer-Cartan set, surjects onto a neighborhood of $0\in\Omega^{1}_{cl,l}$, this then finally implies that $\widetilde{\delta}$ lands in the Maurer-Cartan set (Proposition 2.14). This proves Theorem 2.5.
  6. (6)
    Since $\widetilde{\kappa},\widetilde{\delta}$ are both continuous in the Fréchet topology if $K_{A_{0}}$ is, restricting them to $\mathrm{FC}^{\mathrm{sm}}$ proves Theorem 2.6.
@@ 2.1.2. Formal deformations of flat connections
We recall the following elementary facts about flat connections. If $A_{t}\colon(-\epsilon,\epsilon)\to\Omega^{1}$ is a smooth curve of flat connections, then from differentiating $F_{A_{t}}=0$ we obtain $d_{A_{0}}\dot{A}_{0}=0$, i.e., tangent vectors at $A_{0}$ to curves of flat connections are $d_{A_{0}}$-closed 1-forms. If $\mathrm{g}_{t}\colon(-\epsilon,\epsilon)\to C^{\infty}(M,G)$ is a curve with $\mathrm{g}_{0}\equiv 1$ and $\dot{\mathrm{g}}_{0}=\gamma\in\Omega^{0}(M,\mathfrak{g})$, then the tangent vector at 0 to the curve of flat connections $A_{t}={}^{\mathrm{g}_{t}}A_{0}$ is $\dot{A}_{0}=-d_{A_{0}}\gamma$, i.e. the tangent vector to a curve along the gauge orbit of a flat connection is a $d_{A_{0}}$-exact 1-form. This is equivalent to saying that infinitesimal [^32] By “infinitesimal” we everywhere mean “first-order,” as opposed to formal deformations (of infinite order) discussed below. deformations of flat connections are $d_{A_{0}}$-closed 1-forms, and two such deformations are equivalent whenever they differ by an exact 1-form, i.e. equivalence classes of infinitesimal deformations are in 1-to-1 correspondence with the first twisted cohomology group $H^{1}_{{A_{0}}}$ (sometimes called the Zariski tangent space to the moduli space of flat connections at $A_{0}]$).
@@
The first two equations are satisfied by our assumptions. It is instructive to look at the third equation in detail. We see that we can solve it for $\alpha^{(2)}$ if and only $\frac{1}{2}[\alpha^{(1)},\alpha^{(1)}]$ is $d_{A_{0}}$-exact. Because $\alpha^{(1)}=\alpha$ is closed by assumption, and the bracket is compatible with the differential, $\frac{1}{2}[\alpha,\alpha]$ is always $d_{A_{0}}$ closed. It is exact if and only if $l_{2}^{\prime}(\mathsf{a},\mathsf{a})=0$. In this case, we can write down an explicit solution: